OpenStudy (kingstone):

Find functions f and g so that h(x) = (f ∘ g)(x). h(x) = (6x - 14)^8 A. f(x) = 6x - 14, g(x) = x^8 B. f(x) = 6x^8 - 14, g(x) = -14 C. f(x) = x^8, g(x) = 6x - 14 D. f(x) = (6x)^8, g(x) = -14 I would very much appreciate any kind of help! :)

OpenStudy (anonymous):

you could use \(g(x)=6x-14\) and then \(f(x)=x^8\) that way, \[f\circ g(x)=f(6x-14)=(6x-14)^8\]

OpenStudy (kingstone):

@satellite73 I'm just trying to make sure, so do you believe the answer is C?